OpenStudy (anonymous):

How would you simplify 24√12?

OpenStudy (anonymous):

\[48\sqrt{3}\]

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

the 12 is gotten when multiplying 4 and 3 tto and theres no sqrt of 3 so multiply 24 by 2 and keep the sqrt of 3

OpenStudy (anonymous):

Oh alright, that makes sense now. Thank you!

OpenStudy (anonymous):

no prob!

OpenStudy (anonymous):

\[24\sqrt{2^2 *3}\] \[= 24*2\sqrt3 \] \[=48\sqrt3 \]